OpenStudy (anonymous):

Solve the proportion: 16 over 17 = x over 40 A. 42.5 B. 37.6 C. -42.5 D. -37.6

OpenStudy (anonymous):

b

OpenStudy (anonymous):

16/17 = x/40 16*40=17x 640=17x 640/17=x x=37.6